### v1

# Probing the limits of effective temperature consistency in actively driven systems

Abstract
We investigate the thermodynamic properties of a single inertial probe driven into a nonequilibrium steady-state by random collisions with self-propelled active walkers. The probe and walkers are confined within a gravitational harmonic potential. We evaluate the robustness of the effective temperature concept in this active system by comparing values of distinct, independently motivated definitions: a generalized fluctuation-dissipation relation, a kinetic temperature, and a work fluctuation relation. Our experiments reveal that, under specific conditions, these independent measurements yield a remarkably consistent effective temperature over a wide range of system configurations. Furthermore, we also identify regimes where this consistency breaks down, which delineates the fundamental limits of extending equilibrium-like thermodynamic concepts to athermal, actively driven systems.
I Introduction
A single colloidal particle confined within a harmonic potential and immersed in a thermal fluid offers a fundamental realization of a microscopic statistical thermometer [1, 2]. In thermal equilibrium, the particle’s random displacements follow a Boltzmann (Gaussian) distribution, allowing the thermodynamic temperature, $T$, of the surrounding environment to be directly determined by its mean potential energy via the principle of equipartition [3].
Near and at equilibrium, $T$ can be equivalently defined through various thermodynamic relations, which can be applied to and measured by a colloidal thermometer. Static relations, measured at stationary states, include canonical distributions and energy equipartition, as well as configurational definitions of temperature [4, 5]. Dynamic relations, measured by probing the system’s response properties, are given by the fluctuation-dissipation theorem [6, 7, 8, 9, 10], which rigorously connects the effects of thermal noise and energy dissipation. In addition, fluctuation theorems also relate temperature to fluctuations in thermodynamic quantities such as heat, work, and entropy [11, 12, 13].
Colloidal thermometers have thus been widely employed in experiments, both to validate theoretical measures of equilibrium temperature and to identify deviations from it [14, 15, 16, 17]. Crucially, various definitions of absolute temperature must yield consistent, measurement-independent values in accord with the zeroth law of thermodynamics.
In stark contrast, no universal definition for temperature exists for many natural and engineered systems that operate far from thermal equilibrium. Nonetheless, the notion of effective temperature has frequently arisen as a heuristic extension of equilibrium statistical mechanics to diverse systems operating far from equilibrium, including glassy materials [18, 19, 20, 21], biological suspensions [22, 14, 15, 23, 24, 25, 26], driven granular media [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38, 39, 40, 41], and active matter [42, 43, 44, 45, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55, 17, 56].
In particular, under well-separated timescales, an equilibriumlike fluctuation-dissipation relation (FDR) can become applicable, leading to a definition of an effective temperature, $T_{\text{eff}}$. While the FDR-based $T_{\text{eff}}$ has proven useful in various fields of physics[57, 58, 16, 59, 60], its thermodynamic interpretation and range of applicability remain subjects of ongoing research [61, 62, 63, 64].
A remaining key challenge lies in identifying under which conditions distinct theoretical definitions of effective temperatures yield mutually consistent values far from equilibrium [65, 66]. Although numerous thermodynamic-like criteria have been proposed to establish a meaningful $T_{\text{eff}}$ for nonequilibrium steady-states (NESS), consistency across measurements remains largely supported by empirical evidence [15, 40].
Figure 1: The system:  A. Experimental setup: A Styrofoam ball (diameter $\sim 4$ cm, 11 gr) is trapped in a gravitational harmonic potential, a plastic bowl (diameter 38cm, depth 5cm), and subjected to collisions with $N_{b}=10$ self-propelled bbots (inset: standard bbot, $4\times 1$ cm, $7.1$ gr). The ball is repeatedly perturbed with a uniform air stream created by an external fan along the $x$-axis (white arrow) to test a fluctuation-response relation. To enforce an abrupt onset and release of the perturbation, a mechanical shutter is used (denoted by ’S’).  B. Velocity distributions of bbot assemblies with different $N_{b}=3,4,$ and 10. Instantaneous velocities were extracted from 30 fps recordings of bbot trajectories tracked over 40 minutes.  C. Spatial distribution of 10 bbots within the harmonic trap, with a soft boundary (steep curvature) indicated by a dashed circle of radius $10$ cm.  D, E. Typical 2020 second trajectories of a single bbot in systems with $N_{b}=3$ (D) and $N_{b}=10$ (E) (upper panels), alongside the corresponding tracer trajectories (lower panels).
In this work, a statistical thermometer is realized experimentally, based on a macroscopic tracer particle confined in a harmonic trap and driven into a NESS by random inelastic collisions with self-propelled walkers (see Section 2 and Fig. 1). Previously [67], it was shown that this system satisfies a linear (dynamic) FDR, with a large enough number of walkers. Here, we assess the consistency of the FDR-based $T_{\text{eff}}$ by comparing it to three independent measures: the tracer’s potential energy and modified kinetic energy (at an unperturbed NESS), and a temperature based on a steady-state work fluctuation relation (FR), in a strongly perturbed NESS. Our main result is displayed in Fig. 2: all three independent measurements of $T_{\text{eff}}$ agree across a range of active bath parameters. We further compare our results with theoretical predictions and discuss conditions under which the consistency of $T_{\text{eff}}$ breaks down.
II Experimental setup and basic characterization
[…]
An external airflow can be applied (in the $x$ direction) and abruptly switched-off by a physical shutter, affecting primarily the tracer particle. This mechanism is used to perturb the tracer’s NESS and measure a response function, as detailed in Section 3.1.
Nonequilibrium properties of the active bath.-  In this setting, a single bbot with sufficient inertia typically performs circular motion around the trap center [68], with a preferred chirality (clockwise for Hexbugs). As the number of bbots $N_{b}$ increases, frequent collisions with other bbots (and the tracer) randomize their propulsion direction and result in an active gas-like state (see SM movie 1). In particular, the bbot velocity distribution remains independent of $N_{b}$ whereas its position distribution deviates from the original circular path with large $N_{b}$, as seen in Fig. 1B,C.
Typical bbot trajectories are shown in Fig. 1D,E. For $N_{b}=3$, collisions are rare, and bbots follow long-lived circular paths; for $N_{b}=10$, frequent collisions yield erratic, stochastic motion. Notably, these self-propelled particles exhibit rich dynamics, including alignment with boundaries and forces and emergent collective motion [69]. In contrast, the passive tracer primarily experiences random inelastic collisions with the active bbots (lower panels). These interactions introduce both noise and dissipation, which, in the spirit of the Langevin framework, result in a Brownian-like motion within the harmonic potential. In this work, we focus on the properties of the tracer particle, as detailed below.
Nonequilibrium properties of the tracer.-  The NESS statistics and dynamics of the tracer particle (in unperturbed conditions) are shown in Fig. 3, with the number of bbots $N_{b}$ used as a control parameter. These results were obtained by averaging 375 particle trajectories of 1-minute duration and a time interval of $\Delta t=1/30$ s.
Both the stationary position ($x$) and velocity ($v_{x}$) distributions of the tracer generally deviate from Gaussian behavior. As shown in Fig. 3A, the position distribution evolves with increasing $N_{b}$, transitioning from an exponential form toward a Gaussian-like profile, particularly within the harmonic trap boundary $|x|<10$ cm (see also Ref. [67]). The velocity distributions (Fig. 3B) exhibit an exponential decay for small $N_{b}=3$, with reliable statistics within $|v_{x}|<50$ cm/s. At larger $N_{b}$, the distributions converge to a distinct non-Gaussian shape that is well described by a stretched exponential function.
The stationary dynamics of the tracer are characterized in Figs. 3C,D via the position and velocity autocorrelation functions (ACF). An increase in the rate of dissipative collisions upon adding more bbots, leads to more strongly damped tracer motion. This is reflected in a suppression of oscillations in the relaxation dynamics, and in the reduction of the relaxation time.
Figure 2: Thermometer consistency:  $N_{b}=\{3,4,6,10,12,15,17\}$ bbots. A. Effective temperature measurements obtained using three independent methods: the potential temperature $T_{\text{pot}}=T_{\text{eff}}$ (circles, Eq. 4), the modified kinetic temperature $\tilde{T}_{\text{kin}}$ (squares, Eq. 6), and a constant temperature $T_{\text{FR}}$ derived from a work FR (diamonds, Eq. 7). Notably, these static temperatures validate the FDR of Eq. 2 (for $N_{b}>3$) and define a consistent effective temperature $T_{\text{eff}}$.  B. Exemplary results for three independent steady-states, used for effective temperature measurements under different conditions: (a) an unperturbed state yielding stationary kinetic and potential temperatures ($T_{\text{pot}}$ and $\tilde{T}_{\text{kin}}$); (b) weakly perturbed state characterized by a linear fluctuation-response relation and its associated $T_{\text{eff}}$; and (c) a strongly perturbed state with a biased work distribution yielding $T_{\text{FR}}$ via a fluctuation relation.
The position and velocity ACFs are well fit by the generic solution of a noisy damped harmonic oscillator (Fig. 3E). We focus on the velocity ACF that is given by [2, 70],
$\overline{C}_{vv}(t)=\frac{C_{vv}(t)}{\langle v_{x}^{2}\rangle_{0}}=e^{-t\gamma/2}\left(\cos{\omega t}-\gamma\frac{\sin{\omega t}}{2\omega}\right),$ (1)
where the bar indicates normalization, $\langle v_{x}^{2}\rangle_{0}$ is the (stationary) velocity variance, and $\omega^{2}\equiv\Omega^{2}-\gamma^{2}/4$. Fitting the ACF to Eq. 1 provides system-specific values of the effective timescales: the damping rate, $\gamma$, and the harmonic frequency, $\Omega$.
These timescales are displayed in Fig. 3F as a function of $N_{b}$, showing a transition between underdamped ($\Omega<\gamma/2$) towards critically-damped ($\Omega\approx\gamma/2$) dynamics as $N_{b}$ increases. A reduction in the relaxation time, $2\gamma^{-1}$, corresponds to an increase in collision frequency, evaluated by image analysis as the mean-free time between tracer-bbot collisions $\tau_{c}$ (black line).
At high densities ($N_{b}=15$ and $17$), fast sequential collisions require a higher temporal resolution to capture the tracer’s inertial dynamics. In these cases, the fit parameters are consistent with the critical regime under $\Delta t=1/60$ s, whereas $\Delta t=1/30$ s results in overdamped behavior. These measurements of velocity fluctuations and their corresponding system-specific timescales are further discussed in the context of the kinetic temperature in Sec. 3.2.
III Effective temperatures
In the following sections, we aim to assess whether three independent definitions of effective temperatures can yield mutually consistent values across different conditions and observables. Specifically, we examine the FDR for position and velocity observables under a step-perturbation (3.1),  effective equipartition between kinetic and potential energies (3.2),  and a work FR in a perturbed NESS (3.3).
Figure 3: NESS dynamics and statistics: Data were obtained for active bath configurations with $N_{b}=\{3,4,6,10,12,15,17\}$ bbots, using a tracer of mass $m=1\pm 0.1$ g confined in a potential of stiffness $\kappa=28.2\pm 3$ g/s2. Results present averages over 375 one-minute tracer trajectories recorded at 30 fps for each $N_{b}$.  A. Normalized position ($x$) distributions. Dashed line show fits to exponential ($N_{b}=3$) and Gaussian ($N_{b}=15$) functions.  B. Normalized velocity ($v_{x}$) distributions. An exponential fit is shown for $N_{b}=3$ and a stretched exponential fit for $N_{b}=15$.  C,D. Position (C) and velocity (D) autocorrelation functions (ACFs), showing increasingly damped dynamics with higher $N_{b}$.  E. Position and velocity ACFs for $N_{b}=10$, fitted with Eq. 1 using $\gamma$ and $\Omega$ as fitting parameters, yielding consistent descriptions for both $C_{xx}$ and $C_{vv}$.  F. Extracted relaxation rate ($\gamma$) and trapping frequency ($\Omega$) as functions of $N_{b}$ (upper panel). The evaluated mean-free time between tracer-bbot collisions ($\tau_{c}$) is plotted as frequency versus $N_{b}$ (lower panel).
III.1 FDR temperature
By subjecting the tracer to a small step-perturbation (an abrupt arrest of a constant force $F_{0}$ at $t=0$), a linear FDR can be expressed for both position and velocity observables as [71, 70] (setting $k_{B}=1$):
$R_{x}(t)=\frac{F_{0}}{T_{\text{eff}}}C_{xx}(t),$ (2)
$R_{v}(t)=\frac{F_{0}}{T_{\text{eff}}}C_{vx}(t),$ (3)
where $R_{x}$ and $R_{v}$ are the mean response functions of the position and velocity observables, respectively. The correlation functions in the (long-time) unperturbed NESS are given by a position ACF, $C_{xx}=\langle x(t)x(0)\rangle_{0}$, and a position-velocity cross-correlation function, $C_{vx}=\langle v(t)x(0)\rangle_{0}$. In this context, an effective temperature $T_{\text{eff}}$ is defined as a proportionality constant.
Notably, a static FDR can be recovered in Eq. 2 at $t=0$, with $C_{xx}(0)\rightarrow\langle x^{2}\rangle_{0}$ and $R_{x}(0)\rightarrow\Delta x_{\epsilon}$, where $\langle x^{2}\rangle_{0}$ is the position variance in the unperturbed NESS, and $\Delta x_{\epsilon}$ in the mean displacement. Considering that $F_{0}=\kappa\Delta x_{\epsilon}$, the effective temperature $T_{\text{eff}}$ coincides with the definition of potential energy equipartition,
$T_{\text{pot}}=\kappa\langle x^{2}\rangle_{0}.$ (4)
Thus, implying that $T_{\text{eff}}=T_{\text{pot}}$ can be defined if the dynamic FDR in Eq. 2 is validated, even in the presence of non-Gaussian steady-state distributions.
[…]
Fig. 4C presents the corresponding FDR tests for both position and velocity observables. The system’s full response to the abrupt force arrest at time $t_{1}=t-t_{0}$ is evaluated as $R_{x}(t_{1})=\langle x(t_{1})\rangle-\langle x\rangle_{0}$ and $R_{v}(t_{1})=\langle v(t_{1})\rangle-\langle v\rangle_{0}$, and compared with the unperturbed correlation functions, $C_{xx}$ and $C_{xv}$. Within measurement error, these results validate the FDRs given in Eqs. 2 and 3, with an effective temperature determined by Eq. 4.
Thereby, in this system the tracer’s potential temperature satisfies a dynamic FDR in both position and velocity observables. We note that the signal-to-noise ratio in $R_{v}$ diminishes with increasing $N_{b}$, rendering the velocity response experimentally inaccessible for larger values ($N_{b}>10$). The independent FDR tests of the experimental setups are further provided in the SI.
Fig. 4D presents a parametric plot of the position FDR, showing normalized fluctuation and response quantities across different $N_{b}$ configurations. Specifically, the solid linear line represents validation of Eq. 2 with $T_{\text{eff}}=T_{\text{pot}}$. Violations are mainly observed with the lowest $N_{b}=2$ and $3$, indicating a breakdown of the effective temperature description (as detailed in Ref. [67]). In contrast, with larger $N_{b}$, the tracer’s potential temperature $T_{\text{pot}}$ satisfies a dynamic FDR and is therefore consistent with $T_{\text{eff}}$.
Figure 4: FDR test of the full-response to a step-perturbation:  Results present averages over 375 two-minute step-perturbation sequences (at $t_{0}=60$ s), under weak external airflow applied by a fan at an operating voltage 99 V. A. Example of a typical perturbation sequence $x(t)$, where the fan is turned on for one minute and abruptly turned off for the subsequent minute.  B. Time-dependent mean position $\langle x(t)\rangle$ and mean velocity $\langle v(t)\rangle$, obtained within a two-minute time window.  C. FDR analysis using Eq. 2 (lower panel) and Eq. 3 (upper panel). These results were obtained with $N_{b}=10$.  D. Parametric plot of position FDRs for $N_{b}=\{2,3,4,6,10,12,15,17\}$ realizations, where the effective temperature $T_{\text{eff}}$ determines the linear slope. In equilibrium-like behavior, all FDR data collapse onto a single line with slope $T_{\text{eff}}\sim\kappa\langle x^{2}\rangle_{0}$. Clear deviations from this linearity are observed for the lowest densities ($N_{b}=2$ and 3, triangles), indicating FDR violations.
III.2 Kinetic and potential energy partition
To probe deviations from equilibrium behavior, we naively define the tracer’s kinetic temperature as,
$T_{\text{kin}}=m\langle v^{2}_{x}\rangle_{0}.$ (5)
This definition seems to result in a contradiction between the kinetic and potential temperatures, since deviations from the potential temperature $T_{\text{pot}}$ (Eq. 4) are observed for all $N_{b}$ configurations (solid gray line in Fig. 5A). Namely, $T_{\text{kin}}$ saturates and becomes largely independent of $N_{b}$ in the high-density limit. This clear mismatch is further supported by direct measurements of the effective trapping frequencies $\Omega$, extracted from the tracer’s velocity ACFs (see Fig. 3F). Across all system setups, $\Omega$ yields values lower than the natural frequency $\sqrt{k/m}$, even in an underdamped dynamic regime.
Figure 5: The modified kinetic temperature:  Data are presented for configurations with $N_{b}=\{3,4,6,10,12,15,17\}$ bbots, averaged over 375 one-minute tracer trajectories recorded in unperturbed conditions.  A. Comparison of the standard kinetic temperature $T_{\text{kin}}$ (solid gray line), calculated using the tracer mass $m\approx 1$ g (Eq. 5), the modified kinetic temperature $\tilde{T}_{\text{kin}}$ (solid red line), evaluated using Eq. 6, and the potential temperature $T_{\text{pot}}$ (dashed blue line). Error bars indicate standard deviations. Dot-dashed red lines show deviations in $\tilde{T}_{\text{kin}}$ arising from increasing the measurement time interval $\Delta t$.  B. Separate contributions to $\tilde{T}_{\text{kin}}$: (upper panel) the velocity variance $\sigma_{v}^{2}=\langle v_{x}^{2}\rangle$, and (lower panel) the effective mass $\tilde{m}$, evaluated from stationary ACF dynamics using Eqs. 1 and 6. Gray dot-dashed lines indicate deviations due to increasing $\Delta t$ in both panels.
However, the two effective temperatures agree, once the change in the effective mass is taken into consideration (solid red line in Fig. 5A). We define an effective mass, $\tilde{m}=\kappa\Omega^{-2}$, based on the velocity correlations. Similar mass corrections have appeared in prior studies of particles embedded in active or driven granular media, where persistent athermal fluctuations give rise to an apparent added mass effect [73, 74, 75]. In these cases, mapping the dynamics onto a linear Langevin equation, consistent with the ACF solution in Eq. 1, relies on short persistent times and typically requires a modification of system parameters [76, 77]. Thus, the kinetic temperature definition is modified to,
$\tilde{T}_{\text{kin}}=\tilde{m}\langle v_{x}^{2}\rangle_{0},$ (6)
where $\tilde{m}$ is defined as a $N_{b}$-specific effective mass, under a constant potential stiffness $\kappa$. Deviations of $\tilde{T}_{\text{kin}}$ measurements resulting from low measurement rates are shown as red dashed lines in Fig. 5A.
Fig. 5B shows the separate contributions to $\tilde{T}_{\text{kin}}$: the velocity variance $\langle v_{x}^{2}\rangle_{0}$ (upper panel) and the effective mass $\tilde{m}$ (lower panel), across different $N_{b}$. We note that $\tilde{m}$ deviates from the tracer’s actual mass ($m=1$ g) and results in values that monotonically increase with $N_{b}$. Both measurements show a dependence on the sampling rate ($\Delta t^{-1}$), particularly for large $N_{b}>10$ systems where $\langle v_{x}^{2}\rangle_{0}$ saturates. Nevertheless, under well-resolved velocity fluctuations, the associated modified kinetic temperature $\tilde{T}_{\text{kin}}$ is consistent with the potential temperature $T_{\text{pot}}$, linking tracer dynamics with its linear response through $T_{\text{eff}}=T_{\text{pot}}$.
Figure 6: Steady-state work FR test:  Results present averages over 375 one-minute perturbed tracer trajectories, under a strong external airflow applied by a fan at an operating voltage $13.5$ V. A. Work FR test for $N_{b}=10$. The temperature measurement of $T_{\text{FR}}$, extracted from a fit to Eq. 7, is consistent with the potential temperature ($T_{\text{pot}}$) determined from the unperturbed definition in Eq. 4.  B. Normalized plot of work FR tests across various $N_{b}=\{2,3,4,6,10,12,15,17\}$ configurations. The solid line indicates perfect agreement between the effective temperatures obtained from the work FR and from the equipartition definition, demonstrating thermodynamic-like consistency.
III.3 Effective temperature based on work FR
Here we investigate a work FR in a NESS subjected to a constant force $F_{0}=\kappa\Delta x_{\epsilon}$ (applied at time $t\rightarrow$-$\infty$). An asymmetric work distribution is obtained by the energetic fluctuation around the steady-state, defined as $\delta W(t)=F_{0}\left[x_{\epsilon}(t)-\langle x_{\epsilon}\rangle\right]$. A steady-state work FR can then be expressed as [78, 79, 13],
$\log{\frac{P_{\epsilon}(+\delta W)}{P_{\epsilon}(-\delta W)}}=\frac{\delta W}{2T_{\text{FR}}},$ (7)
where $P_{\epsilon}(\pm\delta W)$ are negative and positive stationary work distributions, and $T_{\text{FR}}$ is defined as an FR-based effective temperature.
[…]
The dashed linear line represents $T_{\text{FR}}=T_{\text{pot}}$, where $T_{\text{pot}}$ is obtained independently in the unperturbed state (a). Namely, the two temperature measures agree, providing an experimental validation of the work FR (Eq. 7). This result can be extended to different $N_{b}$, as shown in Fig. 6B. Here, the solid linear line represents $T_{\text{FR}}=T_{\text{pot}}$. Deviations for large values of the work fluctuations are expected due to under-sampling. Clear violations of Eq. 7 are mainly observed for $N_{b}=3$ and $2$. Thus, for larger $N_{b}$ systems, $T_{\text{FR}}$ is also consistent with the FDR-based measure $T_{\text{eff}}$. Notably, these temperatures were measured in independent measurements under different perturbations.
IV Limitations of a consistent effective temperature
To further assess the range of consistency of effective temperature measures, we repeated the experiments from Section 3 for other system configurations. The standard experimental setup (detailed in Fig. 1 and shown in Fig. 7A) was modified by varying both the passive tracer and the surrounding active particles. Specifically, these modifications included: using a smaller styrofoam tracer particle (Fig. 7B); employing faster active bbots (Fig. 7C); and introducing a heavy tracer particle (Fig. 7D), i.e., heavier than the bbots. Further system details are provided in the SI.
Figs. 7E,F compare the tracer’s modified kinetic temperature ($\tilde{T}_{\text{kin}}$) and potential temperature ($T_{\text{pot}}$) as a function of the number of bbots ($N_{b}$) for these configurations. The solid linear line represents the condition: $\tilde{T}_{\text{kin}}=T_{\text{pot}}$.
In the alternative experimental setups that include a light tracer (Figs. 7B,C), $\tilde{T}_{\text{kin}}$ and $T_{\text{pot}}$ coincide within measurement error and increase with $N_{b}$, consistent with our previous findings (Fig. 5). These systems can also yield consistent $T_{\text{eff}}$ measurements whether via FDR (Eq. 2) or work FR (Eq. 7) tests (see details SI). However, when replacing the lightweight tracers with a heavy tracer, we observe differences between the measured $T_{\text{pot}}$ and $\tilde{T}_{\text{kin}}$ (see Fig. 7F).
Ultra-high-speed videos (SM movie 2) reveal the mechanical origin of this deviation. For lightweight tracers, interactions with the self-propelled bbots occur as discrete impacts, yielding a measurable ballistic regime between collisions. In contrast, heavy tracers can experience rapid, persistent collisions with the bbots that occur on timescales shorter than the experimental sampling interval ($\Delta t$). Upon strong collisions the tracer visibly changes the bbot propulsion dynamics. In addition, the heavy tracer remains embedded among the bbots, whereas a lighter tracer may be kicked upwards, escaping entrapment.
These effects modify the coupling between tracer and environment and may thus lead to the observed discrepancy. We note that the external air-driven perturbation is generally too weak to induce a significant perturbation to the heavy tracer. Due to both effects, the FDR and FR tests fail for all considered $N_{b}$ systems.
Figure 7: Experimental applicability of effective temperature consistency.  A. Standard experimental setup as presented in Fig. 1 and detailed in Sections 2 and 3.  B. Configuration using a smaller Styrofoam tracer particle ($m=0.8$ g, diameter 3.5 cm).  C. Configuration with faster self-propelled bbots (see SI for details).  D. Configuration employing a heavy tracer particle ($m=8.2$ g, diameter 3.8 cm).  E. Comparison of energy partitioning between the modified kinetic temperature $\tilde{T}_{\text{kin}}$ (Eq. 5) and potential temperature $T_{\text{pot}}$ (Eq. 4) as functions of $N_{b}$ for configurations (i), (ii), and (iii).  F. Systematic violations of effective equipartition observed for configuration (D) with the heavy tracer, indicating the breakdown of thermodynamic-like consistency in this regime.
V Summary and discussion
In this work, we have experimentally demonstrated that a passive granular tracer, confined in a harmonic potential and driven by an active fluctuating medium, can sustain a NESS that is well-characterized by a single effective temperature $T_{\text{eff}}$. This effective temperature serves as a unifying parameter that encapsulates both the system’s dynamic response properties and its stationary fluctuations (as detailed in Fig. 2).
We identify three key features that appear essential to this equilibriumlike behavior of effective temperature consistency. First, both the spontaneous fluctuations of the tracer and its response to weak external perturbations are governed by the same physical mechanism, i.e. random collisions with the active medium [67]. This is manifested in the validity of a linear FDR [10, 60].  Second, the light tracer is only weakly coupled to the environment, which minimizes the chance of physical entrapment or caging by bath particles, an effect known to induce long-term memory in the tracer’s dynamics [80, 81].  Finally, the bath dynamics (bbot motion) are largely unaffected by the external driving (airflow) that acts mainly on the tracer particle, mimicking the timescale separation typical to equilibrium Brownian systems [39, 63, 66].
The agreement between different definitions of effective temperature observed here is unexpected for two reasons. First, the noise is a result of non-exponential collision times and the system exhibits non-trivial steady-state distributions, unlike the Poisson shot noise conditions leading to Boltzmann-like steady-states under which such agreement has been predicted [82, 26, 83]. Second, the use of an inertial tracer introduces memory effects in active matter, often breaking Markovian dynamics and precluding a single consistent effective temperature [84, 81, 85, 66]. That this agreement persists despite these deviations from the idealized theoretical framework suggests a broader robustness of effective temperature concepts than previously recognized.
Conflicts of interest
There are no conflicts to declare.
Acknowledgements
DB, RG and YR acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (Grant agreement No. 101002392). BL would like to thank Deutsche Forschungsgemeinschaft for their support (DFG grant LI 1046/10-1).
Supplementary Information
System technicalities
In our experimental setup, the number of bbots ($N_{b}$) is limited due to their tendency to occasionally tumble upon collisions with other bbots within the potential well. This tumbling effect becomes increasingly pronounced at high densities (and high propulsion speeds), leading to the formation of long-lived static clusters at the trap center. Such clustering can prevent the tracer from accessing certain regions in the trap bulk, thus hindering reliable measurements of the its fluctuations and responses. This effect can be partly observed in the stationary (tracer) probability distribution with $N_{b}=17$, as shown in Fig. 8.
In Fig. 9 we present measurements that quantify the chirality of a single bbot in an assembly composed of $N_{b}$ bbots (in the potential well). We evaluate the net chirality of a system by counting the clockwise (+1+1) and anticlockwise (1-1) crossings (across a central line at $x=0$). The results were obtained by tracking 200 bbot trajectories ($t\in[0,5]$ s), and present a sum of $\pm n$ rotations per frame, and a probability distribution $P(n)$. We find that the systems attain a net clockwise chirality that is reduced by increasing $N_{b}$, and therefore the collision frequency. In comparison, the chirality of the passive tracer particle results in small values and vanish for sufficiently high $N_{b}$.
Additional results and measurements
We furthermore provide the following supplemental material:  FDR tests under constant $N_{b}$ and increasing perturbation amplitudes, i.e., fan operating voltage (Fig. 10);  Individual FDR results for position and velocity tracer observables (Figs. 11,12), and work FR results (Fig. 13), with different $N_{b}$. The tracer’s restitution coefficient measurement (Fig. 13); Particularly, these results were obtained with the standard experimental setup, as detailed in the main paper.
Fig. 15 and 16 present FDR and FR tests with alternative system configurations. Namely, we employ faster bbots (15) and a smaller styrofoam ball as a passive tracer (16). The position autocorrelation functions are approximately proportional to the responses with $T_{\text{eff}}=T_{\text{pot}}$, yet the FDR in the velocity observable show violations for short times. The work FR is valid for sufficiently small values, from which a consistent effective temperature can be extracted with $T_{\text{FR}}\approx T_{\text{eff}}$. We note that reliable $T_{\text{eff}}$ measurements across different $N_{b}$ would require additional tuning of the perturbation amplitude and the potential stiffness, to ensure stable dynamics in the system. The FDR tests generally fail for all considered setups that employ a heavy tracer under the air-driven perturbation (Fig. 17).
Figure 8: A.  Snapshot of an experiment with $N_{b}=17$, showing a temporal static cluster of tumbled bbots in the center trap region. B.  The tracer’s position probability distribution for $N_{b}=17$, showing a flat region in the center, reflecting the physical exclusion by static bbot clusters.
Figure 9: The system’s chirality. Typical single bbot and tracer trajectories, in a system with $N_{b}=3$ (A) and $N_{b}=10$ (B). C.  The distribution of bbot instantaneous rotations $\pm n$ across a central line at $x=0$, where $n>0$ indicates favored clockwise direction. Obtained by tracking 200 bbots trajectories ($t\in[0,5]s$). The upper lines are the means of $P(n)$, showing that net chirality is reduced by increasing $N_{b}$. D.  The same chirality analysis for the passive tracer.
Figure 10: The FDR test under different perturbation amplitudes. A.  The mean displacement $\Delta x$ under a perturbation of a tracer in a $N_{b}=10$ bbot bath, as a function of fan operating voltage $V$ (error bars are standard errors). The applied force $F_{0}=k\Delta x$ increases above V=11V=11V. The tracer’s position standard deviation, $\sigma_{x}$, is used to define the range of small perturbations (solid line). The dashed line is a linear fit of the first 4 data points, with $\Delta x=0.19\cdot V$, whereas V=13V=13V and 13.513.5V are out of the linear regime. B.  The generalized FDRs with $T_{\text{eff}}\sim\langle\Delta x^{2}\rangle_{0}$ are plotted for V=9, 12, 13.5V=9,\ 12,\ 13.5 V, with $N_{b}=10$. The results present an average over $M=375$ perturbation sequences.
Figure 11: Position ($x$) FDR results with different numbers of bbots.  A,B. FDRs with $T_{\text{eff}}\sim\langle\Delta x^{2}\rangle_{0}$ for $N_{b}=\{2,3,4,6,10,12,15,17\}$ bbot baths. The results were obtained for an ensemble average of $M=375$ perturbation sequences, with the same tracer $m\approx 1$ g, gravitational stiffness $k\approx 28.2~\text{g/s}^{-2}$, and fan operating voltage (99 V) with $\Delta x_{\epsilon}<\sigma_{x}$. The vertical dashed lines are $T_{c}=1$ s.
Figure 12: Velocity ($v_{x}$) FDR results with different numbers of bbots.  A,B. FDRs with $T_{\text{eff}}\sim\langle\Delta x^{2}\rangle_{0}$ for $N_{b}=\{2,3,4,6,10,12,15,17\}$ bbot baths. Same details as in Fig. 14.
[…]
Figure 14: Tracer’s coefficient of restitution versus impact velocities. The restitution coefficient $r$ was determined from the sound signal emitted by the styrofoam ball (the tracer) bouncing repeatedly off a surface, as detailed in Ref. [86]. We obtain $r$ by measuring the time lag between consecutive impacts: $r_{i}=|v^{\prime}_{i}/v_{i}|=(t_{i+1}-t_{i})/(t_{i}-t_{i-1})$, for an impact velocity $v_{i}=\frac{g}{2}(t_{i-1}-t_{i})$, $i=2,3,$.... Here, $v_{i}$ and $v^{\prime}_{i}$ are the normal pre- and post-collision velocities occurring at time $t_{i}$, and $g=9.81\textrm{m/s}^{2}$ is the gravitational acceleration. The plot shows four successive impacts of a repeated experiment (50 trials), with mean values from last to first respectively: $r=0.79$ (circle), $0.78$ (diamond), $0.76$ (square), $0.74$ (triangle), for $v_{\perp}=68.37$, $87.35$, $113.4$, $152.83$ cm/s.
Figure 15: Fast bbot setup.  A. The active bath contains $N_{b}=6$ fast bbots, the experimental setup is otherwise the same as in Figs. 10-13.  B. Average individual speed of a single bbot walking in a parabolic arena. Standard bbots (circle) and fast bbots (diamond), as referred to in the main paper.  C. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 11,12. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$.  D. Work FR test as detailed in Fig. 13.
Figure 16: Small tracer setup.  A. A small styrofoam ball is used as a passive tracer (33 cm diameter, $0.8$ g). The active bath contains $N_{b}=10$ standard bbots.  B. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 11,12. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$.  C. Work FR test as detailed in Fig. 13.
Figure 17: Heavy tracer setup.  A. A heavy plastic ball is used as a passive tracer (33 cm diameter, $8.2$ g). The active bath contains $N_{b}=10$ standard bbots.  B. FDR test for the tracer’s position and velocity observables, under the same perturbation setup as detailed in Figs. 11,12. A strong airflow ($13.5$ V fan operating voltage) was used as a step-perturbation. The effective temperature is $T_{\text{eff}}=\kappa\langle x^{2}\rangle_{0}$. The correlation functions fail to capture the response to the step-perturbation.